### v2

# False Discovery Rate and Localizing Power

Abstract
False discovery rate (FDR) is commonly used for correction for multiple testing in neuroimaging studies. However, when using two-tailed tests, making directional inferences about the results can lead to vastly inflated error rate, even approaching 100% in some cases. This happens because FDR control the error rate only globally, over all tests, not within subsets, such as among those in only one or another direction. Here we consider and evaluate different strategies for FDR control with two-tailed tests, using both synthetic and real imaging data. Approaches that separate the tests by direction of the hypothesis test, or by the direction of the resulting test statistic, more properly control the directional error rate and preserve FDR benefits, albeit with a doubled risk of errors under complete absence of signal. Strategies that combine tests in both directions, or that use simple two-tailed p-values, can lead to invalid directional conclusions, even if these tests remain globally valid. To enable valid thresholding for directional inference, we suggest that imaging software should allow the possibility that the user sets asymmetrical thresholds for the two sides of the statistical map. While FDR continues to be a valid, powerful procedure for multiple testing correction, care is needed when making directional inferences for two-tailed tests, or more broadly, when making any localized inference.
keywords:
False discovery rate, two-tailed tests, localizing power, subsetting property
[…]
1 Introduction
In the context of the multiple testing problem, control over the false discovery rate (fdr) provides an alternative to the control over the familywise error rate (fwer): while fwer refers to the chance of even one type i (false positive) error is committed in a set (family) of statistical tests, fdr refers to the proportion of such errors among tests for which the null hypothesis has been rejected, that is, the proportion of false discoveries among all discoveries. These topics have been extensively covered in the literature; for an early review of the multiple testing problem in brain imaging, see Nichols and Hayasaka (2003); for a review in genetics, also broadly applicable to imaging, see Goeman and Solari (2014); for early papers and for an overview of the development of fdr, see Seeger (1968); Sorić (1989); Benjamini and Hochberg (2000); for its introduction to neuroimaging, see Genovese et al. (2002); for a critique, see Chumbley and Friston (2009); for a discussion of various fdr approaches, see Korthauer et al. (2019). Methods for fwer correction in brain imaging are often based on the random field theory (rftWorsley et al., 1996, 2004), permutation tests (Holmes et al., 1996; Nichols and Holmes, 2002; Winkler et al., 2014), Monte Carlo simulations (Poline and Mazoyer, 1993; Forman et al., 1995; Cox, 1996) and occasionally, the method based on Bonferroni (1936).
Two hallmarks of most neuroimaging studies are that they can involve thousands of statistical tests, and that one can reasonably expect effects to be observed in both directions for many of these. Controlling the fdr aims to provide more power than fwer in the setting where many null hypotheses are false. This benefit has seen fdr widely adopted in neuroimaging. However, the behavior of fdr when signal manifests in positive direction for some tests, and negative for others, has not received as much attention, even if it is a pertinent problem in the field. In principle, using fdr with two-tailed tests is straightforward: the two-tailed p-values are obtained, then an fdr-controlling procedure is applied. As long as inference remains focused on all tests, using fdr on two-tailed p-values has nothing special compared to using one-tailed p-values. Using two-tailed tests can become problematic if, after doing the test, the researcher proceeds to draw inferences on each direction: fdr is not controlled within one direction or another, only globally. Previous literature on directional hypotheses pertinent to this problem can be found in Shaffer (1995, 2002); Williams et al. (1999); Benjamini and Yekutieli (2005). The same problem emerges if a researcher attempts to draw inferences within regions of a map of p-values after applying fdr globally across all of them. Both problems are a general manifestation of the fact that fdr, while a powerful procedure, does not have the ability to allow localization within subsets of tests, that is, it lacks localizing power. As one of our reviewers pointed out, “making discoveries in either direction while controlling the fdr is trickier than it may seem”.
In this note we start by briefly revisiting the multiple testing problem in imaging, p-value adjustment, the use of fdr, and discuss interpretation for two-tailed tests in the light of statistical properties of fdr-controlling procedures. We demonstrate that merely applying fdr does not control the error rate on each side of the statistical map when the researcher wishes to make inferences about the direction after using two-tailed tests, and can lead to invalid conclusions when effects are asymmetrically distributed across the two sides. We investigate methods that can be considered to address such issues, with focus on a slight modification of the method proposed by Benjamini and Yekutieli (2005). We also bring attention to the fdr approach proposed by Benjamini et al. (2006), which is more powerful than the one originally proposed by Benjamini and Hochberg (1995), and which has seldom been used in brain imaging. We also offer a method for p-value adjustment using this more powerful procedure. While the paper is presented in a neuroimaging context, the same topics are relevant to other fields in which fdr is widely used.
2 Theory
2.1 Preliminaries
Consider a number $V$ of voxels (or vertices, or edges, or faces) of a representation (image) of the brain. For each voxel, a null hypothesis $\mathcal{H}^{0}_{i}$, $i\in\{i,\ldots,V\}$ is tested through the computation of some test statistic $t_{i}$, and a corresponding (one-tailed) p-value $p_{i}$ that represents the probability of finding $t_{i}$ at least as large as the one observed. Each p-value can be considered statistically significant if equal or below some pre-defined test level $\alpha$ (typically 0.05). If the distribution of the test statistic is symmetric around zero and continuous, two-tailed p-values can be obtained from one-tailed as $p^{\text{two}}=1-\left|2p^{\text{one}}-1\right|$. More generally, if the distribution is either continuous or discrete, two-tailed p-values can be obtained as $p^{\text{two}}=2\cdot\min\left\{p^{\text{one}},1-p^{\text{one}}+C\right\}$, where $C=0$ for continuous (often parametric) p-values, or $C=1/J$ for discrete, permutation-based (non-parametric) p-values, and $J$ is the number of permutations. If the distribution of the test statistics is asymmetric, two-tailed p-values must be obtained from that distribution, even if one-tailed p-values are available.
2.2 The multiple testing problem
[…]
Regardless of the degree of dependency among tests, it may be the case that, faced with the task of screening many such tests, a researcher may be willing to tolerate a small amount of them to be incorrectly declared significant (false discoveries), as long as a great majority of tests are correctly declared as such (true discoveries), and the errors are limited; if no discoveries are made, there are no false positives by definition. In other words, the researcher may be willing to accept some false discoveries as long as their proportion over all discoveries does not spill over some pre-defined threshold $q$ (often $q$ == $0.05$). Controlling this proportion is the essence of fdr approaches.
2.3 The basic fdr procedure: bh
Control over the fdr can be achieved with the procedure by Benjamini and Hochberg (1995) (bh): let $p_{(1)}\leqslant\ldots\leqslant p_{(V)}$ be the ordered p-values. Find the largest index $i$ for which:
$p_{(i)}\leqslant\frac{iq}{V}$ (1)
where $q$ is the upper limit on the proportion of false discoveries that the researcher is willing to tolerate. Then reject all null hypotheses that have p-values smaller than or equal to that critical $p_{(i)}$. Graphically, one would lay the ordered p-values in a Cartesian plane that has as horizontal axis the index $i$; draw a line with slope $q/V$ passing through the origin, then reject all tests whose ordered p-values lie to the left of the highest p-value at or below that line (Figure 1, left panel).
[…]
When the tests are independent, the bh procedure controls the fdr at the level $qV_{0}/V$, where $V_{0}$ is the (unknown) number of true null hypotheses. However, complete independence is not strictly necessary: Benjamini and Yekutieli (2001) showed that it is sufficient that tests exhibit a form of dependence that they termed positive regression dependency on subsets (prds). In this form of dependence, knowledge that one p-value where the null is true is small does not make less likely that the other (V1)(V-1) p-values are also small. If this condition is satisfied, the bh procedure still controls the fdr. Analyses where a false positive does not make the other tests less likely to be significant, as typical in brain imaging, satisfy the prds condition.222prds itself is defined in Benjamini and Yekutieli (2001); additional literature on forms of dependency between variables, relevant to prds, can be found in Lehmann (1966); Esary et al. (1967); Sarkar (1969); Kimeldorf and Sampson (1989).
A procedure that controls the fdr more generally, for any dependence between tests (as opposed to only under independence or prds) was proposed by Benjamini and Yekutieli (2001) (by). The procedure is similar to bh, except for a small modification in Equation 1, which becomes $p_{(i)}\leqslant iq/V/c(V)$, where $c(V)=\sum^{V}_{j=1}1/j$. Graphically, this results in a less inclined threshold line, and thus, a smaller cutoff $p_{(i)}$ (more conservative) than with the original bh procedure.
2.4 A more powerful fdr procedure: bky
Unfortunately, bh and by are both conservative in the sense that the expected fdr is below $qV_{0}/V$, which is always less than or equal to the desired level $q$. To mitigate this conservativeness, Benjamini et al. (2006) proposed to estimate $V_{0}$ after a first pass using bh with $q^{\prime}=q/(1+q)$, then repeat with $q^{*}=q^{\prime}V/\hat{V}_{0}$ if at least one hypothesis was rejected in the first pass; this two-pass procedure has been proposed for use in brain imaging (Chen et al., 2009), albeit it remains rarely used. In their original paper, Benjamini et al. (2006) recognised that there is no reason to stop after the second pass; in fact, the procedure can be repeated iteratively until there are no more rejections. Doing so leads to following procedure (bky), which we adopt here: starting from the lowest p-value, keep rejecting hypotheses for the successive $p_{(i)}$ as long as at least one $p_{(j)}$, $j\geqslant i$, satisfies:
$p_{(j)}\leqslant\frac{jq}{V+1-i(1-q)}$ (2)
stopping at the first $i$ for which the inequality is no longer satisfied for any $j\geqslant i$; all subsequent nulls, inclusive, are retained (Benjamini et al., 2006, Definition 7).
As with bh, bky can also be interpreted graphically (Figure 1, right panel): for each ordered p-value $p_{(i)}$, trace a threshold line with slope q/(V+1i(1q))q/(V+1-i(1-q)) passing through the origin, and check if any p-value from that position (inclusive) onward is at or below that line; if so, reject the corresponding null hypothesis, and move to the next p-value, now with a new threshold line that is steeper (more liberal) than the previous. The bky procedure differs from bh, however, not only on the use of an adaptive threshold line, whose slope varies according to the position $i$, but also on the stopping criterion: whereas bh uses as cutoff the largest $p_{(i)}$ that satisfies its inequality, bky stops at the first $p_{(i)}$ that no longer satisfies it. This has importance when computing adjusted p-values, discussed next.
2.5 Adjustment of p-values
It is convenient to inspect a map of p-values that incorporates correction for multiple testing: thresholding such map at the desired test level immediately reveals significant results and is straightforward to examine in most imaging software via, e.g., a slider, or by varying the upper or lower limits of the color scale, or by changing image contrast window center and width. In contradistinction, maps of uncorrected p-values need be thresholded at some corrected $\alpha$ or $q$ levels. Corrected p-values for the bh procedure can be computed by finding the smallest $q$ over every position $i$ that still satisfies the inequality given in Equation 1:
$q_{(i)}^{\textsc{bh}}\geqslant\frac{p_{(i)}V}{i}=\left[p_{(i)}^{\textsc{bh}}\right]_{\text{corr}}$ (3)
where the inequality becomes equality at the lowest (strictest) value of $q_{(i)}$ for which the respective $p_{(i)}$ is significant (the notation with brackets $[$ ]] avoids double subscripts). In a similar manner, for bky, corrected p-values can be found by solving Equation 2 for $q$ at $i$:
$q_{(i)}^{\textsc{bky}}\geqslant\min_{j\geqslant i}\left\{\frac{p_{(i)}(V+1-i)}{j-ip_{(i)}}\right\}=\left[p_{(i)}^{\textsc{bky}}\right]_{\text{corr}}$ (4)
Unfortunately, corrected p-values for bh and bky are not guaranteed to be monotonically related to their corresponding uncorrected p-values (that is, corrected p-values not necessarily increase or decrease as their corresponding uncorrected p-values increase or decrease). Graphically, the corrected p-values correspond to the slope of the line that connects each point to the center of the Cartesian grid, and that slope can be higher than $q$ even for p-values considered significant. Such lack of monotonicity is problematic because thresholding a map of such corrected p-values at some desired level $q$ will not produce the same result as thresholding the map of original p-values at the critical $p_{(i)}$ identified with the respective fdr-controlling method. One may expect monotonicity when doing statistical tests: it allows one to trade-off between maximizing power while minimizing false positives (Zeisel et al., 2011). Following Yekutieli and Benjamini (1999, Definition 2.4), adjusted p-values, which replace corrected p-values without having this issue, can be computed for the bh procedure as:
$\left[p_{(i)}^{\textsc{bh}}\right]_{\text{adj}}=\min_{j\geqslant i}\left\{\left[p_{(j)}^{\textsc{bh}}\right]_{\text{corr}}\right\}$ (5)
which can be read as the cumulative minimum of the corrected p-values, starting from the largest. Because bky uses a different stopping criterion than bh, the adjustment is also different: rather than using the smallest corrected p-value from the current position $i$ onward, bky requires the largest corrected p-value up to the current position $i$, that is, the cumulative maximum:
$\left[p_{(i)}^{\textsc{bky}}\right]_{\text{adj}}=\max_{j\leqslant i}\left\{\left[p_{(j)}^{\textsc{bky}}\right]_{\text{corr}}\right\}$ (6)
[…]
With adjusted p-values, significance thresholds no longer depend on the position $i$ for bky: maps of adjusted p-values can be arbitrarily thresholded at any $q$ in the interval between 0 and 1 (Figure 2). It should be noted that ties are very common with adjusted p-values; they pose no problem for interpretation. Interestingly, using Equation 6 for bky, as opposed to a formulation equivalent to Equation 5, solves the issue with the paradoxical results observed by Reiss et al. (2012), whereby adjusted p-values would seem smaller than their unadjusted counterparts.
[…]
2.6 Localizing power
Methods for fwer and fdr control differ not only on the measurement of error that is controlled, but on at least another aspect: if a procedure controls the fdr at some level for a set of hypotheses, this control does not necessarily extend to every subset of hypotheses, and the actual fdr can in fact be much higher (or even much lower) for any subset; a procedure that controls the fwer for a set of hypotheses, however, also controls the fwer for every subset. This feature, that fwer procedures have, but fdr procedures have not, was described by Finner and Roters (2001), and called subsetting property by Goeman and Solari (2014). In the context of imaging, we find the term localizing power (or the lack of it) preferable as it immediately conveys the meaning that further localization within parts of an image may or may not be possible, depending on the procedures used for inference (Friston et al., 1996; Nichols and Hayasaka, 2003; Noble et al., 2022).
The ability to localize effects within parts of an image requires that the error rate is controlled within that part. Not having localizing power is problematic, for example, if the researcher wishes to draw conclusions within specific regions after fdr has been applied globally. Likewise, this becomes problematic if the researcher wishes to draw conclusions pertaining to the direction of effects — that is, inferring about each tail separately — after applying fdr on two-tailed p-values. To illustrate that, consider these two hypothetical examples:
Example 1: Direction error
An investigation into the association between cortical thickness (measured at every vertex of a surface representation of the brain) and a behavioral measure involves a sample of hundreds of individuals. Acknowledging that effects could be manifest in either direction, and pursuing statistical rigor, the researcher employs two-tailed p-values, with adjustments for multiple testing using an fdr. The map of test statistics is then masked at the significant level $q=0.01$, thus retaining for visualization only the significant vertices. The analysis uncovers large areas in the frontal and parietal lobes with a positive relationship to the behavioral measure, as hypothesized. However, it also reveals a negative association spanning broad portions of the fusiform gyrus, as well as vertices showing positive or negative associations scattered across the cortex. The researcher conducting the analysis writes a report detailing the novel findings in the negative direction, emphasizing the stringent methodologies that incorporated two-tailed tests and a strict fdr significance threshold, well below the conventional level $q=0.05$, and concludes (erroneously) that only about 1% of the negative associations are expected to be false positives.
Example 2: Localization error
In a study of a compound for reducing the severity of relapses in multiple sclerosis (ms), white matter hyperintensities (wmh) are analyzed voxelwise. Subjects are randomly allocated into active compound and placebo groups. Based on prior knowledge and on random population sampling, normality is assumed and a two-sided group comparison is conducted using a two-sample $t$-test. This produces two-tailed p-values calculated from the Student’s $t$-distribution. Control for multiple testing across voxels is achieved using fdr at $q=0.05$. After determining the fdr threshold, the inverse cumulative distribution function of the Student’s $t$-distribution is used to derive the critical threshold $t_{\textsc{fdr}}$. Given the two-tailed nature of the test, symmetrical thresholds $t_{\textsc{fdr}}$ and $t_{\textsc{fdr}}$ are set in an image visualization tool. The results show a notable reduction in the periventricular wmh load in the group that used the active compound, as expected (and hoped), as well as mixed results in various juxtacortical regions, albeit with mostly increases in wmh load in these areas. In the report, the investigator concludes (erroneously) that the drug may amplify the wmh load in juxtacortical areas given that, as fdr was used, no more than 5% of the observed changes are false positives.
At first, it may seem that the researchers in both examples did everything correctly, going as far as using two-tailed tests and ensuring that the multiplicity of tests would be addressed rigorously. A casual reader of the respective study reports may find no issue. Yet, where it comes to statements about the direction or localization of effects, the interpretation is not that, among the findings in the negative direction (Example 1) or in a particular region (Example 2), the proportion of false positives is within $q$, but only that the proportion of false positives among all tests that were subjected to fdr is within $q$. The actual, realized (but unknown) fdr is not guaranteed to be within 1% for the negative (or positive) direction in Example 1, nor 5% within the juxtacortical white matter (or any other region) in Example 2, and can in both cases be anywhere between 0 and 100% in the reported direction or region.
These issues are a consequence of the lack of the subsetting property. Localization errors can be avoided by refraining from making statements or taking actions that could refer to the amount of errors in specific regions when fdr is controlled only globally, or by doing separate fdr analyses within each subset (Efron, 2008). The same holds to avoid direction errors. Direction errors can also be avoided, we assert, by partitioning the statistical map into two disjoint sets: one with the positive test statistics, another with the negative test statistics, then proceeding to fdr correction in each subset using their two-tailed p-values. This is expected to work because under the null hypothesis, the partitioning using the sign of the test statistic is independent of the two-tailed p-values; the latter are distributed uniformly between 0 and 1 and, in either side of the map, more extreme test statistics (stronger effects) imply lower p-values. Thus, the requirements for fdr, using either bh or bky, are satisfied. The fdr can then be controlled at the desired level $q$ for the positive and for the negative sides of the map. We verify this in the next section, along with other methods.
3 Evaluation
3.1 Synthetic data
To assess the realized (empirical) fdr across procedures we used synthetic data formed by 2000 tests, which could represent voxels in a group-level result image. For each, a test statistic $z$ was simulated following a normal distribution with zero mean and unit variance. Five different scenarios were considered for independent test statistics: (i) complete null, with no synthetic effects added; (ii) positive effects only, in which 25% of the tests had their $z$ statistic increased by 3; (iii) negative effects only, in which 25% of the tests had their $z$ statistic decreased by 3; (iv) both positive and negative effects, balanced, in which 25% of the tests had their $z$ statistic increased by 3, another 25% had their $z$ statistic decreased by 3; and (v) both positive and negative effects, unbalanced, in which 10% of the tests had their $z$ statistic increased by 3, another 40% had their $z$ statistic decreased by 3. Another five scenarios (vix) were considered; these were identical to the first five except that instead of independent tests, the 2000 test statistics were made non-independent, with a correlation of 0.25. In each scenario, uncorrected p-values were computed by referring to the normal cumulative distribution function. A summary of the scenarios is shown in Table 1.
Table 1: Scenarios used to evaluate the different fdr approaches with synthetic data. In each scenario, 2000 test statistics were simulated and subjected to fdr procedures. The process was repeated 2000 times. Scenarios iv contain independent tests; tests in scenarios vix have nonzero dependency, as indicated by the “Correlation” column.
Simulation scenario Tests with positive effects Tests with negative effects Correlation among tests
i 0% 0% 0
$i$ 25% 0% 0
iii 0% 25% 0
iv 25% 25% 0
v 10% 40% 0
vi 0% 0% 0.25
vii 25% 0% 0.25
viii 0% 25% 0.25
ix 25% 25% 0.25
x 10% 40% 0.25
Simulations for these 10 scenarios were repeated (realized) 2000 times. Each was assessed with both bh and bky. For this simulated data evaluation, we use $q=0.05$. Four different strategies were considered, described below and summarized in Table 2. Hence, we have 80 sets (10 scenarios, 2 fdr procedures, and 4 strategies) of results, each based on 2000 realizations.
Table 2: Summary of the four strategies considered for running fdr.
Strategy Inputs to fdr Calls to fdr
Canonical $\{p^{\text{one}}_{i}\}$, $\{1-p^{\text{one}}_{i}\}$ Two, each with $V$ tests
Combined $\{p^{\text{one}}_{i},1-p^{\text{one}}_{i}\}$ One, with $2V$ tests
Two-tailed $\{p^{\text{two}}_{i}\}$ One, with $V$ tests
Split-tails $\{p^{\text{two}}_{i}:z_{i}>0\}$, $\{p^{\text{two}}_{i}:z_{i}<0\}$ Two, with a total of $V$ tests
For each synthetic voxel $i\in\{1,2,\ldots,V\}$, $p^{\text{one}}_{i}$ refers to their one-tailed p-values, $p^{\text{two}}_{i}$ refers to their two-tailed p-values; $z_{i}$ are their corresponding test statistics (for example, it could be the result of a signed statistical test, such as the $t$-test).
  • Canonical, in which adjustment is applied to one-tailed p-values in one direction, and to one-tailed p-values in the opposite direction, that is, applied separately to $p$ and to 1p1-p. This procedure produces two complete maps of adjusted p-values, one for each direction; alternatively, it produces two thresholds, one for each direction.
  • Combined, in which adjustment is applied simultaneously to one-tailed p-values in the two directions, i.e., on the set formed by $p$ and 1p1-p. This procedure produces two complete maps of adjusted p-values, one for each direction; alternatively, it produces one threshold, to be applied to both directions.
  • Two-tailed, in which two-tailed p-values are subjected to adjustment. This procedure produces one complete map of adjusted p-values; alternatively, it produces one threshold.
  • Split-tails, in which the tests are separated into two disjoint subsets, one with positive ($z>0$) and another with negative ($z<0$) statistics; then the two-tailed p-values within each subset are subjected to adjustment. For the cases in which the test statistic is zero ($z=0$), the uncorrected two-tailed p-value is 1, and so is the adjusted p-value; such pin risk is extremely unlikely with continuous test statistics. After the two disjoint sets are joined, this procedure produces one complete map of adjusted p-values; alternatively, it produces two thresholds, one for positive test statistics, one for negative test statistics.
Counting true and false positives
For each realization, true and false positives are considered in relation to the true signal; for directional inferences (positive or negative), these refer to the results in the positive and negative sides of the statistical map; for non-directional (both sides), the directions of the test statistic and of the true signal are ignored; these definitions are formalized in Table 3. The proportion of false discoveries within a realization is the number of false discoveries divided by the total number of discoveries.
Assessing power
The power of each of the fdr methods and the four different strategies is the ratio between the number of true discoveries by the number of tests in which true signal was added, computed for reach realization. For directional inferences (positive or negative), these refer to the results in the positive and negative sides of the statistical map; for non-directional (both sides), the directions of the test statistic and of the true signal are ignored; as described in Table 3. Thus, power here represents a distinct quantity in relation to the usual statistical power. Here it represents the sensitivity of a multiple testing procedure judged by its ability to recover a number of tests that contain true signal, that is, tests in which the null hypothesis is false.
Table 3: Counting of tests that meet criteria for containing a true synthetic effect in a given direction, for having the sign of the test statistic matching the direction of the signal, and being statistically significant after an fdr-controlling procedure. These counts are used to calculate the empirical proportion of false discoveries and empirical power with synthetic data.
Direction Counts
Both sides:
[…]
True discoveries $\{i:(p^{\text{adj}}_{i}\leqslant q)\text{ and }(\top^{+}_{i}\text{ or }\top^{-}_{i})\}$
False discoveries $\{i:(p^{\text{adj}}_{i}\leqslant q)\text{ and }\text{not }(\top^{+}_{i}\text{ or }\top^{-}_{i})\}$
Positive side:
[…]
Negative side:
True effects $\{i:\top^{-}_{i}\}|$
[…]
For each synthetic voxel $\in\{1,2,\ldots,V\}, i+\top^{+}_{i}$ and $\top^{-}_{i}$ indicate whether a true positive or true negative effect was added to that voxel, $z_{i}$ represents the corresponding test statistic at that voxel, and $p^{\text{adj}}_{i}\leqslant q$ represents a significant adjusted p-value at that voxel after an fdr procedure. For each realization, the counts are the sizes (cardinalities) of each of the sets. The empirical fdr within a realization is the number of false discoveries observed in a given direction divided by the total number of discoveries in that direction, whereas power is the number of true discoveries in a given direction divided by the number of tests with true effects in that direction. For computations that refer to both sides, the direction is ignored. Averaging the proportion of false discoveries or the calculated power across realizations yields the results shown in Tables 4 and 5, respectively.
The overall empirical fdr (observed false discovery proportion) and power were computed by averaging across realizations. Confidence intervals (95%) were calculated with the Wald method (Brown et al., 2001).
3.2 Real data
To assess these four different strategies for two-tailed fdr with real data, as well as to provide an empirical comparison between bh and bky, we use a statistical map originally produced as part of the Neuroimaging Analysis Replication and Prediction Study (narpsBotvinik-Nezer et al., 2020). The narps was an effort to assess the variability of results among different analysis teams analyzing the same imaging dataset. In the study, the teams were provided with the same raw functional magnetic resonance imaging (fmri) data, collected during a gambling task. Each team analyzed the data to investigate predefined hypotheses about brain activation related to gains and losses, and employed diverse analysis workflows, with no two teams choosing identical pipelines. We use (for no particular reason) the maps produced by team 9U7M for narps Hypothesis 1. Note that the specifics of the hypothesis are not relevant to the present analysis; consult Botvinik-Nezer et al. (2020) for details about data acquisition, the predefined hypotheses, the overall narps methodology, and its data provenance. For this real data evaluation, we use $q=0.05$.
4 Results
4.1 Synthetic data
Tables 4 and 5 summarize the results for the analyses using synthetic data across the 10 different scenarios that involved varying proportions of true positive and negative effects, as well as dependencies among tests. For a given test, the ideal result is that the confidence interval includes the chosen $q=0.05$ value (represented in the Table in percentage, i.e., 5%).
Table 4: (page 4.1) Comparison of the false discovery proportions (%) among different fdr methods and correction strategies, when false positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The nominal proportion of false discoveries is $q\cdot 100\%=5\%$; their corresponding confidence intervals (95%) are shown between parentheses.
Table 5: (page 4.1) Comparison of power (%) among different fdr methods and correction strategies, when true positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The nominal proportion of false discoveries is $q\cdot 100\%=5\%$; their corresponding confidence intervals (95%) are shown between parentheses.
Benjamini–Hochberg Benjamini–Krieger–Yekutieli
Scenario Canonical Combined Two-tailed Split-tails Canonical Combined Two-tailed Split-tails
Both sides:
i 9.5 (8.2–10.8) 5.1 (4.1–6.0) 5.1 (4.1–6.0) 9.6 (8.3–10.8) 9.5 (8.2–10.8) 4.8 (3.8–5.7) 4.8 (3.8–5.7) 9.6 (8.3–10.8)
$i$ 3.8 (3.7–3.8) 3.7 (3.7–3.8) 3.7 (3.7–3.8) 3.0 (3.0–3.1) 4.7 (4.6–4.7) 4.1 (4.1–4.2) 4.5 (4.5–4.6) 4.3 (4.3–4.4)
iii 4.0 (3.9–4.0) 4.0 (3.9–4.0) 4.0 (3.9–4.0) 3.2 (3.2–3.3) 4.9 (4.9–4.9) 4.2 (4.2–4.3) 4.7 (4.6–4.7) 4.6 (4.5–4.6)
iv 2.6 (2.6–2.6) 2.6 (2.6–2.6) 2.6 (2.6–2.6) 2.6 (2.6–2.6) 3.2 (3.2–3.2) 3.2 (3.2–3.2) 4.3 (4.3–4.3) 4.3 (4.3–4.3)
v 2.6 (2.6–2.6) 2.6 (2.6–2.6) 2.6 (2.6–2.6) 2.3 (2.3–2.3) 3.7 (3.7–3.7) 3.1 (3.1–3.2) 4.3 (4.2–4.3) 4.2 (4.2–4.3)
vi 7.7 (6.5–8.9) 4.5 (3.5–5.4) 4.5 (3.5–5.4) 5.4 (4.4–6.4) 8.1 (6.9–9.2) 4.0 (3.2–4.9) 4.0 (3.2–4.9) 4.7 (3.8–5.6)
vii 3.9 (3.6–4.3) 3.8 (3.5–4.1) 3.8 (3.5–4.1) 2.5 (2.3–2.7) 5.3 (4.9–5.8) 4.2 (3.9–4.6) 4.6 (4.3–5.0) 3.7 (3.3–4.0)
viii 4.4 (4.1–4.8) 4.2 (3.9–4.5) 4.2 (3.9–4.5) 2.9 (2.6–3.1) 6.0 (5.6–6.5) 4.6 (4.3–4.9) 5.0 (4.7–5.3) 4.0 (3.7–4.3)
ix 3.7 (3.5–3.9) 2.7 (2.5–2.8) 2.7 (2.5–2.8) 2.9 (2.7–3.0) 4.8 (4.5–5.1) 3.2 (3.1–3.4) 4.3 (4.1–4.5) 4.6 (4.3–4.8)
x 3.3 (3.1–3.4) 2.5 (2.4–2.7) 2.5 (2.4–2.7) 2.4 (2.3–2.6) 4.9 (4.6–5.2) 3.2 (3.0–3.3) 4.3 (4.1–4.4) 4.3 (4.1–4.5)
Positive side:
i 5.3 (4.4–6.3) 2.9 (2.2–3.7) 2.9 (2.2–3.7) 5.4 (4.4–6.4) 5.3 (4.4–6.3) 2.9 (2.1–3.6) 2.9 (2.1–3.6) 5.4 (4.4–6.4)
$i$ 3.8 (3.7–3.8) 1.9 (1.9–1.9) 1.9 (1.9–1.9) 3.0 (3.0–3.0) 4.7 (4.6–4.7) 2.1 (2.1–2.1) 2.3 (2.3–2.4) 4.3 (4.3–4.4)
iii 4.8 (3.8–5.7) 99.7 (99.4–99.9) 99.7 (99.4–99.9) 5.9 (4.9–7.0) 3.5 (2.7–4.3) 99.9 (99.8–100.0) 100.0 (99.9–100.0) 5.0 (4.0–5.9)
iv 2.5 (2.5–2.6) 2.5 (2.5–2.5) 2.5 (2.5–2.5) 2.5 (2.5–2.6) 3.1 (3.1–3.1) 3.1 (3.1–3.1) 4.2 (4.2–4.2) 4.2 (4.2–4.2)
v 2.5 (2.5–2.6) 6.0 (5.9–6.1) 6.0 (5.9–6.1) 3.6 (3.5–3.7) 2.7 (2.6–2.7) 7.3 (7.2–7.4) 9.8 (9.7–9.9) 4.4 (4.3–4.5)
vi 4.3 (3.5–5.2) 2.6 (1.9–3.4) 2.6 (1.9–3.4) 3.0 (2.3–3.8) 3.9 (3.1–4.7) 2.1 (1.5–2.8) 2.1 (1.5–2.8) 2.4 (1.7–3.0)
vii 3.6 (3.3–3.9) 1.8 (1.7–2.0) 1.8 (1.7–2.0) 2.3 (2.2–2.5) 4.7 (4.3–5.0) 1.9 (1.7–2.1) 2.2 (2.0–2.4) 3.2 (3.0–3.4)
viii 3.6 (2.8–4.4) 70.1 (68.1–72.1) 70.1 (68.1–72.1) 3.1 (2.4–3.9) 3.8 (2.9–4.6) 70.5 (68.5–72.5) 74.6 (72.7–76.5) 2.8 (2.1–3.5)
ix 2.4 (2.2–2.5) 1.9 (1.8–2.0) 1.9 (1.8–2.0) 2.0 (1.9–2.1) 3.1 (2.9–3.4) 2.4 (2.3–2.6) 3.4 (3.2–3.5) 3.4 (3.2–3.6)
x 2.2 (2.0–2.4) 3.6 (3.4–3.8) 3.6 (3.4–3.8) 2.4 (2.3–2.6) 2.7 (2.4–2.9) 4.7 (4.4–4.9) 6.3 (6.0–6.6) 3.3 (3.1–3.5)
Negative side:
i 4.5 (3.6–5.5) 2.5 (1.8–3.2) 2.5 (1.8–3.2) 4.5 (3.6–5.5) 4.3 (3.5–5.2) 2.1 (1.5–2.7) 2.1 (1.5–2.7) 4.3 (3.5–5.2)
$i$ 4.2 (3.3–5.1) 99.8 (99.5–100.0) 99.8 (99.5–100.0) 5.8 (4.8–6.8) 4.0 (3.1–4.9) 99.9 (99.8–100.0) 100.0 (99.9–100.0) 5.7 (4.7–6.7)
iii 4.0 (3.9–4.0) 2.1 (2.1–2.2) 2.1 (2.1–2.2) 3.2 (3.2–3.3) 4.9 (4.8–4.9) 2.3 (2.2–2.3) 2.5 (2.5–2.5) 4.6 (4.5–4.6)
iv 2.7 (2.7–2.7) 2.7 (2.6–2.7) 2.7 (2.6–2.7) 2.7 (2.6–2.7) 3.3 (3.2–3.3) 3.3 (3.2–3.3) 4.4 (4.3–4.4) 4.4 (4.3–4.4)
v 2.6 (2.6–2.6) 1.7 (1.7–1.7) 1.7 (1.7–1.7) 2.0 (2.0–2.0) 3.9 (3.9–3.9) 2.0 (2.0–2.1) 2.8 (2.7–2.8) 4.2 (4.2–4.2)
vi 3.4 (2.6–4.1) 1.8 (1.2–2.4) 1.8 (1.2–2.4) 2.4 (1.7–3.0) 4.2 (3.3–5.0) 1.9 (1.3–2.5) 1.9 (1.3–2.5) 2.4 (1.7–3.0)
vii 2.8 (2.1–3.5) 70.4 (68.4–72.4) 70.4 (68.4–72.4) 2.4 (1.7–3.0) 3.0 (2.3–3.8) 73.5 (71.6–75.4) 76.3 (74.4–78.2) 2.9 (2.2–3.6)
viii 3.7 (3.5–4.0) 2.0 (1.8–2.1) 2.0 (1.8–2.1) 2.5 (2.3–2.6) 5.3 (5.0–5.7) 2.3 (2.2–2.5) 2.7 (2.5–2.9) 3.7 (3.5–4.0)
ix 2.7 (2.5–2.9) 2.2 (2.0–2.3) 2.2 (2.0–2.3) 2.3 (2.1–2.4) 3.5 (3.3–3.8) 2.7 (2.5–2.8) 3.7 (3.4–3.9) 3.8 (3.5–4.0)
x 2.6 (2.4–2.7) 1.5 (1.4–1.7) 1.5 (1.4–1.7) 1.8 (1.6–1.9) 4.1 (3.9–4.4) 1.9 (1.8–2.0) 2.7 (2.5–2.9) 3.6 (3.4–3.8)
Benjamini–Hochberg Benjamini–Krieger–Yekutieli
Scenario Canonical Combined Two-tailed Split-tails Canonical Combined Two-tailed Split-tails
Both sides:
i
$i$ 77.4 (77.3–77.5) 66.5 (66.3–66.6) 66.5 (66.3–66.6) 73.3 (73.2–73.4) 81.6 (81.5–81.7) 68.4 (68.2–68.5) 70.3 (70.2–70.4) 80.2 (80.0–80.3)
iii 77.7 (77.6–77.8) 66.8 (66.7–66.9) 66.8 (66.7–66.9) 73.6 (73.5–73.7) 81.7 (81.6–81.8) 68.4 (68.3–68.5) 70.4 (70.3–70.6) 80.3 (80.2–80.4)
iv 76.4 (76.3–76.5) 76.4 (76.3–76.5) 76.4 (76.3–76.5) 76.4 (76.3–76.5) 80.0 (79.9–80.0) 80.0 (79.9–80.0) 85.3 (85.2–85.3) 85.3 (85.2–85.3)
v 78.8 (78.7–78.8) 76.4 (76.3–76.5) 76.4 (76.3–76.5) 76.8 (76.8–76.9) 84.1 (84.0–84.2) 80.0 (79.9–80.1) 85.3 (85.2–85.4) 87.0 (86.9–87.1)
vi
vii 76.4 (75.3–77.6) 64.2 (63.1–65.4) 64.2 (63.1–65.4) 73.2 (72.3–74.2) 80.4 (79.1–81.7) 65.1 (63.9–66.2) 66.8 (65.6–68.0) 79.6 (78.7–80.5)
viii 76.5 (75.4–77.7) 64.4 (63.3–65.6) 64.4 (63.3–65.6) 73.2 (72.3–74.2) 81.8 (80.4–83.2) 66.5 (65.3–67.7) 68.3 (67.0–69.5) 80.6 (79.6–81.5)
ix 74.8 (74.6–75.0) 76.4 (76.3–76.6) 76.4 (76.3–76.6) 75.9 (75.8–76.1) 78.4 (78.2–78.6) 80.0 (79.8–80.1) 85.2 (85.1–85.3) 85.0 (84.9–85.2)
x 77.4 (76.9–77.9) 75.7 (75.2–76.3) 75.7 (75.2–76.3) 76.3 (75.8–76.8) 82.6 (82.0–83.2) 79.0 (78.5–79.6) 84.2 (83.6–84.7) 86.4 (86.0–86.9)
Positive side:
i
$i$ 77.5 (77.4–77.6) 65.3 (65.2–65.4) 65.3 (65.2–65.4) 73.4 (73.3–73.5) 81.7 (81.6–81.8) 67.0 (66.9–67.2) 68.8 (68.7–68.9) 80.3 (80.1–80.4)
iii
iv 76.5 (76.4–76.6) 76.5 (76.4–76.6) 76.5 (76.4–76.6) 76.5 (76.4–76.6) 80.0 (79.9–80.1) 80.0 (79.9–80.1) 85.3 (85.2–85.4) 85.3 (85.2–85.4)
v 62.1 (61.9–62.3) 79.3 (79.1–79.4) 79.3 (79.1–79.4) 68.8 (68.6–69.0) 63.1 (62.9–63.3) 83.6 (83.5–83.8) 90.7 (90.5–90.8) 72.9 (72.7–73.1)
vi
vii 76.4 (75.3–77.6) 63.8 (62.6–64.9) 63.8 (62.6–64.9) 73.3 (72.3–74.2) 80.1 (78.7–81.4) 64.4 (63.2–65.6) 66.1 (64.9–67.3) 79.4 (78.5–80.4)
viii
ix 74.6 (73.5–75.7) 76.3 (75.5–77.1) 76.3 (75.5–77.1) 75.7 (74.8–76.6) 78.0 (76.9–79.1) 79.7 (78.9–80.6) 85.0 (84.2–85.8) 84.6 (83.8–85.5)
x 60.3 (59.0–61.6) 78.3 (77.5–79.2) 78.3 (77.5–79.2) 68.2 (67.2–69.2) 61.9 (60.6–63.2) 82.8 (82.0–83.6) 89.2 (88.4–90.0) 73.0 (71.9–74.0)
Negative side:
i
$i$
iii 77.8 (77.7–77.9) 65.7 (65.5–65.8) 65.7 (65.5–65.8) 73.7 (73.6–73.8) 81.8 (81.7–82.0) 67.1 (67.0–67.2) 68.9 (68.8–69.1) 80.4 (80.3–80.5)
iv 76.5 (76.4–76.7) 76.5 (76.4–76.6) 76.5 (76.4–76.6) 76.5 (76.4–76.6) 80.2 (80.0–80.3) 80.2 (80.1–80.3) 85.5 (85.4–85.6) 85.5 (85.4–85.6)
v 83.1 (83.0–83.1) 75.8 (75.7–75.9) 75.8 (75.7–75.9) 79.0 (78.9–79.0) 89.5 (89.4–89.6) 79.2 (79.1–79.3) 84.1 (84.0–84.2) 90.7 (90.6–90.7)
vi
vii
viii 76.5 (75.3–77.6) 63.9 (62.8–65.1) 63.9 (62.8–65.1) 73.2 (72.3–74.2) 81.7 (80.3–83.1) 65.9 (64.7–67.2) 67.7 (66.4–68.9) 80.6 (79.6–81.5)
ix 75.1 (74.0–76.2) 76.7 (75.9–77.5) 76.7 (75.9–77.5) 76.3 (75.3–77.2) 78.9 (77.7–80.0) 80.3 (79.5–81.1) 85.6 (84.8–86.4) 85.6 (84.7–86.4)
x 81.8 (80.8–82.7) 75.2 (74.3–76.0) 75.2 (74.3–76.0) 78.4 (77.5–79.2) 87.9 (86.9–88.9) 78.2 (77.3–79.1) 83.0 (82.1–83.9) 89.9 (89.1–90.7)
4.1.1 Correction strategies
In the absence of signal, that is, under the complete null hypothesis, and with independent test statistics (scenario i), the canonical method worked well when false positives were counted among observed positive test statistics only, or among observed negative test statistics only. In these cases, the empirical false discovery proportion was within the confidence intervals, being neither conservative nor anti-conservative. When false positives were counted among all test statistics (both sides, positive and negative), the observed error rate was twice as large as for each side separately, and thus, twice as large as the nominal test level. Similar behavior was observed for the split-tails method. Combined and two-tailed methods controlled the error rate at the nominal level, albeit became conservative when only one or the other direction was considered.
With only positive or only negative true signal (scenarios ii and iii), all four correction strategies produced error rates that were within the expected level for test statistics with the same sign as the signal, even if leaning heavily towards conservativeness. However, for test statistics with the opposite sign as the signal, the observed proportion of false discoveries was not controlled by a large margin, approaching 100% error rate. When both sides were considered to count false positives (that is, all tests considered), the error rate was controlled within the nominal level, leaning again towards conservativeness, albeit less heavily.
With both positive and negative true signal in a balanced scenario (scenario iv), the error rate was controlled regardless of how the error rate was counted (positive test statistics only, negative only, or both). That control, however, leaned towards conservativeness. With unbalanced amounts of signal (scenario v), which more typically represents real experiments, the error rate was not controlled in the side opposite to the one with a preponderance of true signal, with the nominal error rate lying outside the confidence interval, and conservative on the opposite side.
4.1.2 fdr methods
While the same patterns above were observed for both bh and bky methods, bky had the advantage of leading to observed false discovery rates that were closer to the nominal level in the presence of signal with the same direction as the observed test statistic, as opposed to leaning towards conservatism as bh. In other words, whenever there was signal, bky was more powerful than bh. However, this also meant that the problems with excess false discoveries observed in the side opposite from where there was more signal was also exacerbated.
4.1.3 Dependencies among tests
The same patterns above were observed for dependent tests (scenarios vix). Introduced dependencies did not substantially alter the results when comparing the correction strategies or the different fdr methods.
4.2 Real data
Figure 3 shows maps thresholded at positive and negative sides using the different fdr methods and correction strategies, as well as with uncorrected p-values; the thresholds in each case ($t$ statistics) are also shown.
Figure 3: Thresholded statistical maps using different fdr methods and correction strategies for directional inference, as well as uncorrected, for the narps data. Significant results in the positive side are those greater than the threshold $t_{\text{pos}}$; in the negative side, those smaller (more negative) than $t_{\text{neg}}$; if nothing is declared significant, the threshold is set, for display purposes, at a value more extreme than the test statistic at that side, here generically represented as $\infty$. The canonical and split-tails strategies produce generally similar results; likewise, combined and two-tailed methods produce broadly similar maps. Inferences using combined or two-tailed maps should not make directional statements about the results; these are only possible in the canonical and split-tails methods. The figure also shows slight but consistent gains in power with the bky procedure when compared to the usual bh procedure. Only visualization software that allow asymmetrical thresholds can be used to display fdr-corrected results that enable directional inferences.
The uncorrected map is thresholded, at the positive side, at $t_{\text{pos}}=1-t_{\text{cdf}}^{-1}(\alpha/2,\nu)=+1.982$, where $t_{\text{cdf}}^{-1}$ is the inverse cumulative distribution function of the Student’s $t$-distribution, $\nu=107$ are the degrees of freedom for this analysis, and $\alpha=0.05$ is the test level; it is thresholded similarly at the negative side at $t_{\text{neg}}=t_{\text{cdf}}^{-1}(\alpha/2,\nu)=-1.982$. The uncorrected thresholded maps suggests the presence of signal in both positive and negative directions, with a preponderance of effects in the negative side, involving more brain regions, particularly around the cortex.
Similar computations were done for each of the correction methods in Figure 3, which shows the thresholds. The canonical strategy produced no significant results in the positive side of the map; the same for the split-tails strategy. The combined strategy resulted in tests considered significant with both positive and negative test statistics. Since all p-values (for tests investigating positive and negative effects) are pooled in the correction, the thresholds are symmetrical. Here, fewer and smaller regions are identified as significant on the negative side (where there is a preponderance of signal) and some results on the positive side that were not seen with the canonical strategy appear significant. The two-tailed strategy produced broadly similar results as with the combined method. Finally, the split-tails strategy produced results that were visually generally similar as the canonical fdr. While the differences were minute in relation to bh, bky consistently produced slightly lower thresholds, leading to a slight increase in power.
5 Discussion
5.1 Two-tailed tests
Two-tailed tests trace their roots to the first decades of statistics as a scientific discipline. Fisher (1925, p. 47) noted that “in some cases we wish to know the probability that the deviation, known to be positive, shall exceed an observed value, whereas in other cases the probability required is that a deviation, which is equally frequently positive and negative, shall exceed an observed value”. This allows assessment of extreme outcomes in any direction, not restricting hypotheses to a specific one. Similarly, in brain imaging, there are cases in which interest lies on only one of the tails, whereas in others, the interest lies on both tails of the distribution of test statistics. Consider, for example, an fmri study where the same information is presented via visual or auditory cues: some brain areas may be selectively modulated by vision and others by hearing; interesting comparisons involve both directions, i.e., higher involvement of some regions during visual than auditory, and vice-versa. Alternatively, consider a voxel-based morphometry study that compares gray matter volume between professional singers and professional pianists: again one might expect brain regions show differences in one group while other regions show the reverse pattern.
In effect, using two-tailed tests as default has at times been recommended for neuroimaging studies: Chen et al. (2019) argued that “it is difficult to imagine a scenario where this researcher would test all voxels simultaneously for just one directional change”. However, the researcher must be aware that the amount of errors is not guaranteed within a given direction if fdr is applied across all tests. Merely applying fdr methods as integrated into neuroimaging packages such as afni, fsl, FreeSurfer, or spm — each referenced in thousands of peer-reviewed publications annually (Poldrack et al., 2019) — and then proceeding to directional inferences, does not control the error rate within each direction, as illustrated by the synthetic and real data.
5.2 Control of the error rate
The analyses using synthetic data indicate that, while fdr-based adjustment (with either bh or bky) controls the error rate in the set of tests to which the procedure is applied, should the researcher decide to make inferences about direction of effects, extra care is necessary, otherwise such inferences can be affected by excess error rate, which can well amount to up to 100% of the tests about which inference is made if true signal is present in only one side. Simple strategies such as using a two-tailed test followed by fdr, or combining the two sets of results with the directional hypotheses into one single fdr correction, do not guarantee control over the error rate for either direction, even though these control the error rate globally, i.e., when all tests are considered. Such global test, however, provides no information on the direction of effects.
Alternative approaches were considered here. For example, running fdr separately for the two set of p-values with each of the directional hypotheses (the canonical strategy), or running fdr separately for the positive and negative test statistics (the split-tails strategy). The analyses demonstrate that these approaches are able to control the error rate in each direction. However, they come with an important penalty: should the null hypothesis be true everywhere, the error rate, which should be at the level $q$ under the complete null across both sides, is doubled, approaching $2q$ when both sides are considered. Should the researcher be interested in retaining the error controlled even in the complete null, then the test level should be halved. It should be noted that trying to control fdr for all subsets simultaneously would transform it into an fwer-controlling procedure (Finner and Roters, 2001).
While the simulations using synthetic data allow quantification of the observed proportion of false discoveries and how it diverges from the nominal, expected rate, real data allows for the visualization of the different patterns that may be observed in typical studies, depending on the fdr method and correction strategy adopted. With real data, the canonical and split-tails lead to broadly similar statistically significant spatial maps; likewise, the combined and two-tailed strategies were broadly similar one to another. Two-tailed tests can be seen as a particular case of multiple testing correction, since they consider both directions of an effect. Hence, it is not surprising that the the two-tailed strategy produces nearly identical results as the combined strategy, which is similar to correction across contrasts (Alberton et al., 2020).
It should be understood, however, that areas of positive signal that appeared as significant with the two-tailed and combined strategies in the narps dataset example may have emerged simply because there is a preponderance of signal in the negative side of the map; evidence for that comes from the results using synthetic data. This is because procedures that control fdr are adaptive, that is, they start with a stringent control for the smallest p-value, and proceed in steps, each contingent on the findings of the previous one, progressively loosening stringency and continuing as long as each subsequent test continue to be declared significant (Benjamini and Hochberg, 2000; Nichols, 2007). With two-tailed and combined strategies, the large amount of tests with strong signal in one side of the map (negative in our example) allows more opportunities for tests in the opposite side (positive) to be declared significant. The canonical and split-tails strategies eliminate the opportunity for eventual strong or widespread signals present in one side impact the other, as fdr is applied separately to each side.
5.3 Localizing power
Even though the above issues are a direct consequence of fdr not possessing the subsetting property, i.e., not providing localizing power, routine strategies for data analysis and visualization may obscure this property. For example, the use of symmetrical thresholds is common in imaging visualization software: the user loads a statistical map and applies a threshold $t$; values above +t+t are shown in a hot color scheme (e.g., red to yellow), while values below t-t are shown in a cool color scheme (e.g., blue to cyan); for two-tailed tests, this seems appropriate and rigorous. Yet, the amount of errors in any direction is not controlled with fdr, only the global amount of errors, something that is not always apparent to the user of such tools, or they may (incorrectly) assume that the fdr is controlled in the positive and negative sides of the map. Another situation is when drawing conclusions about the amount of errors within a region after controlling globally; it might seem that controlling the error rate for the whole brain is the most rigorous, such that inferences within sub-regions would be conservative, given that a larger number of tests is taken into account. Yet, the amount of errors within any particular region is not controlled, only the global amount of errors. Any attempt to isolate a particular direction or a particular subset of tests will fail to guarantee the desired proportion of false discoveries within that subset.
5.4 More power with bky
An important yet little appreciated fact is that the bh procedure — the most commonly adopted — does not control the amount of false discoveries at the exact level $q$ desired by the researcher, but at a lower level $qV_{0}/V$ that is determined by the (unknown) proportion of tests in which the null hypothesis is true among the total number of tests (Section 2.3). Hence, the procedure is, in fact, conservative in relation to the level desired by the researcher. Despite this, it remains more powerful than procedures that control fwer such as Šidák or Bonferroni that use, for all tests, a fixed significance threshold that determines whether a p-value is significant; unless $V_{0}=V$ (complete null), $\textsc{fwer}=1$ for any fdr-controlling procedure. Whereas bh is adaptive in the sense that the threshold varies according to the sorted p-value at position $i$, with a fixed slope given by $q/V$, the bky procedure is even more adaptive than bh in that the slope itself varies according to the sorted p-value at position $i$ as q/(V+1i(1q))q/(V+1-i(1-q)) (Figure 1). The quantity q/(V+1i(1q))q/(V+1-i(1-q)) is only smaller than $q/V$ if i<1/(1q)i<1/(1-q). Thus, for typical $q$, usually well below 0.5, bky is more powerful than bh as long as smallest ($i=1$) test is rejected. As far as power considerations are concerned, the bky procedure is therefore preferable over the bh procedure.
This extra power comes with a computational price: bh runs in $\mathcal{O}(V)$ time, whereas bky runs in $\mathcal{O}(V^{2})$ time. This means that, while the time it takes to run fdr grows linearly with the number of tests to be corrected, for bky it grows quadratically. Another point worth mentioning is that bky-adjusted p-values can easily be above 1 (Figure 2). This is not a unique feature of bky: other multiple testing procedures can also lead to adjusted p-values being higher than 1, the simplest being Bonferroni, that leads to adjusted p-values that are equal to the uncorrected p-values multiplied by the number of tests $V$. This fact does not have substantial implications for hypotheses testing, and can be solved via truncation at 1.
The bky procedure was developed as an improvement over bh, and thus it would seem that the same considerations regarding dependencies between tests would hold. However, results have only been proven for the independent case (Benjamini et al., 2006), even if it tends to perform well in simulations under certain forms of dependence (e.g., Stevens et al., 2017). Scaling the thresholds of bky by $1/c(V)$ to accommodate arbitrary dependencies, as in by, while compelling, has not been proven to control the fdr either (nor that it does not).
5.5 Dependencies among tests
While we only considered one case of dependencies among tests, and observed adequate control over the error rate, existing results in the literature (Genovese et al., 2002; Logan and Rowe, 2004; Reiner-Benaim, 2007) demonstrate that, for typical cases, the modification proposed by Benjamini and Yekutieli (2001) for a more robust control over the false discovery proportion (the by procedure) is not necessary, and that even in the presence of negative correlations, bh, being more substantially more powerful, is preferable over by. That extends to bky as well.
5.6 Visualization of results
As mentioned in Section 2.6 and 5.3, localization and direction errors can be avoided by refraining from making localized or directional statements encompassing subsets of the tests, making only such statements when they refer to the whole set of tests that were subjected to the fdr procedure. Direction errors can also be avoided by using strategies such as those that we named canonical or split-tails. These approaches, however, require that the user is able to display the statistical maps in software that enables non-symmetric thresholds to be used. That is, the ideal imaging tool would accommodate separate thresholds for positive and for the negative sides of the map, such that these can be set individually after running fdr with one of these strategies. This can work also with approaches that do not use abrupt thresholds, but instead a transparency gradient, allowing subthreshold evidence be viewed informatively while highlighting suprathreshold results (Allen et al., 2012; Taylor et al., 2025).
Currently, very few software packages allow the user to easily set asymmetric thresholds for positive and negative sides of the map. Having this feature available across different imaging software would more easily allow for directional investigations with fdr and visualization of results, even using two-tailed tests. For example, a map of uncorrected p-values can be thresholded at the fdr-controlling level, or a map of adjusted p-values can be thresholded at the desired level. Alternatively, if the distribution of the test statistic is known, the threshold p-value can be converted to a threshold statistic (using the cumulative distribution function, parametric or non-parametric), which can be applied to partition statistical map into significant and non-significant regions. A thresholded map (of statistics or of p-values) can be binarized and used as a mask to display or select regions for further analyses; the same workaround can be used for multiple regions, each of them subjected to correction separately, possibly with the test level further scaled down by the number of such regions should the researcher wish to control the error rate under the complete null.
5.7 Recommendations
When applying the fdr in neuroimaging studies with two-tailed tests, researchers should adapt their approach to their goals. fdr does not ensure error control within subsets, so localized inferences require applying fdr separately, possibly with the test level adjusted to accommodate the case of no effects anywhere. For directional inferences, such as detecting positive or negative effects, the canonical or split-tails strategies are recommended to control the fdr within each direction, though the total fdr may double in the complete absence of effects; the test level is proportionally reduced if that is a concern. To apply this approach, software have to allow thresholding to be asymmetric. For non-directional inferences, the two-tailed or combined strategies suffice. The bky procedure offers greater power than the bh method.
6 Conclusion
While fdr procedures provide more power than fwer methods if many tests are expected to be significant and the user is willing to accept some false positives, caution is warranted when making directional inferences after using two-tailed tests. Strategies like running fdr separately on positive and negative test statistics, or on one-tailed p-values partitioned by hypothesis direction, avoid direction errors and control the false discovery proportion in each tail. However, these strategies lose fdr benefits under the complete null, such as controlling the fwer in these cases. Researchers should refrain from making directional inferences about fdr-corrected maps after two-tailed tests particularly when signals are not symmetrically distributed, since weak control means the proportion of false discoveries is only guaranteed globally; for the same reason, localization of effects within subsets is not possible using these methods. The bky procedure allows for additional power, albeit it does not eliminate issues with asymmetric signals. Allowing asymmetric thresholds in neuroimaging software would enable appropriate visualization after partitioned fdr.
Source code
Code related to this paper is available at https://github.com/andersonwinkler/fdrfun.
Author contributions
All authors contributed to all aspects of this research.
Competing interests
None.
Acknowledgements
Support was provided by the National Institutes of Health. A.M.W.: U54-HG013247; P.A.T.: ZIC-MH002888; T.E.N.: R01-EB026859, U24-DA041123, and U19-AG073585; C.R.: RF1-MH133701 and P50-DC014664. The authors are grateful to the anonymous reviewers who provided helpful feedback.